What is the future value of $7,890 at the end of 5 periods at 8% compounded interest?

What is the present value of $7,890 due 9 periods hence, discounted at 6%?

What is the future value of 16 periodic payments of $7,890 each made at the end of each period and compounded at 10%?

What is the present value of $7,890 to be received at the end of each of 19 periods, discounted at 5% compound interest?
Answer #1
1) Future Value = Present value * (1+i)^n Where,
= 7890*(1+0.08)^5 i = 8%
= $     11,593.00 n = 5
2) Present value = Future value * (1+i)^-n Where,
= 7890*(1+0.06)^-9 i = 6%
= $        4,670.08 n = 9
3) Future Value = Annual cash flow * Future value of annuity of 1
= $        7,890.00 * 35.94973
= $ 2,83,643.37
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.10)^16)-1)/0.10 i = 10%
= 35.94972986 n = 16
4) Present Value = Annual cash flow * Present value of annuity of 1
= $        7,890.00 * 12.08532
= $     95,353.18
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.05)^-19)/0.05 i = 5%
= 12.08532086 n = 19